OpenStudy (zyberg):

How to solve: \(\sqrt{x - 2} - \sqrt{x + 10} = 2\)? (The answer is that there are no solutions, however, I can't seem to work it out...)

OpenStudy (3mar):

Are you familiar with logarithmic functions?

OpenStudy (zyberg):

No, I am not.

OpenStudy (chareye18):

u first need to square everything

OpenStudy (3mar):

2 mins plz

OpenStudy (phi):

maybe one way is to do this: \[ \sqrt{x - 2} - \sqrt{x + 10} = 2 \\ \sqrt{x - 2} = 2+\sqrt{x + 10} \] now square both sides to get \[ x-2 = 4+x+10 +4 \sqrt{x+10} \] after we simplify we get \[ - 4= \sqrt{x+10} \] and (because we only use the "principal square root" i.e. the positive root) there is no solution.

OpenStudy (zyberg):

Thank you very much!